# UFO3: Weaving the Digital Agent Galaxy [Uncaptioned image]

### v1

@@ 1 Introduction
The rise of intelligent agents wang2024survey marks a new era of human–computer interaction, where large language models (LLMs) naveed2025comprehensive are evolving from text-based reasoning engines qu2025tool into autonomous digital operators capable of perceiving, acting, and coordinating across tasks zhang2024large. Yet despite this progress, most agent frameworks remain confined within a single device or platform, be it a browser tab ning2025survey; zheng2024gpt, a desktop environment zhang2025ufo; zhang2025ufo2, or a mobile app zhang2025appagent; wang2024mobile. This confinement sharply limits their ability to harness the rich, distributed computational ecosystem that modern users inhabit.
When an agent is trapped within one operating system, it cannot access the complementary strengths of other devices, such as GPU clusters for computation, desktop applications for document editing, or mobile sensors for context capture. The result is a fragmented landscape of intelligent but siloed agents: each powerful in isolation yet collectively underutilized. This gap between reasoning ability and real-world actuation leaves vast potential untapped. To truly advance the next frontier of automation and reasoning, agents must operate beyond the boundaries of any single device or OS, forming a coherent digital collective where Windows laptops, Linux servers, mobile devices, and edge nodes collaborate seamlessly houben2017opportunities; brudy2019cross to gather and act upon ubiquitous intelligence.
@@
  • Declarative decomposition into a dynamic DAG (TaskConstellation). Natural-language or programmatic requests are decomposed by the global ConstellationAgent into a structured DAG bei2025graphs of TaskStars and TaskStarLines that encode workflow logic and dependencies. This declarative structure is amenable to automated scheduling, introspection, and rewriting throughout execution.
  • Continuous, result-driven graph evolution. The TaskConstellation is a living data structure. Intermediate outputs, transient failures, and new observations trigger controlled rewrites, adding diagnostic TaskStars, creating fallbacks, rewiring dependencies, or pruning completed nodes, so the system adapts dynamically instead of aborting on errors wu2024agentkit.
@@
  • Unified Agent Interaction Protocol (AIP). Built atop persistent WebSocket channels, we develop AIP, a protocol that provides a unified, secure, and fault-tolerant layer for agent registry, session management, task dispatch, and coordination. It ensures reliability under network fluctuations through automatic reconnection and retry, while exposing a lightweight, extensible interface that allows new agents to integrate seamlessly into the UFO3 ecosystem yang2025survey.
  • Template-driven framework for MCP-empowered device agents. To democratize agent creation, UFO3 provides a lightweight development template and toolkit for rapidly building new device agents. Developers can declare capabilities, bind to local environments, and extend them through one or more Model Context Protocol (MCP) servers hou2025model for tool augmentation. This modular design accelerates integration while maintaining consistency across the constellation.
@@
In essence, UFO3 dissolves device boundaries and transforms the digital estate into a single, adaptive collaborator brudy2018investigating; marks2020multi; zhang2018towards. It unifies distributed devices into a cohesive digital organism, one that executes user intents safely, asynchronously, and efficiently across heterogeneous environments. At its core, the Agent Interaction Protocol (AIP) serves as the connective tissue of this ecosystem, roviding a unified, fault-tolerant, and extensible communication substrate that allows new agents to join seamlessly and interoperate reliably. Over time, multiple constellations can interconnect through AIP, weaving together agents, devices, and capabilities into a self-organizing Digital Agent Galaxy. Through this design, UFO3 redefines cross-device automation, elevating it from a brittle engineering challenge to a unified orchestration paradigm, where multi-device workflows become naturally expressive and scalable across the landscape of ubiquitous computing.
@@ 2.1 Digital Agents
Leveraging the power of LLMs, modern digital agents have emerged as powerful interfaces bridging human intent and the complex software ecosystems people interact with daily zhang2024large. These agents can parse natural language requests, interpret screenshots and system state zheng2025vem; wu2025gui; zhao2025learning, decompose complex goals into subtasks huang2024understanding, and generate scripts or commands to execute tasks using a variety of tools wang2024survey. Execution modalities include API calls, operating-system-level commands, GUI interactions via automation or accessibility interfaces, and code generation zhang2025api; qiao2023taskweaver; wang2024large; zhang2025swe.
Digital agents operate across a wide spectrum of platforms. Web-based agents can navigate browsers and search the Internet ning2025survey; zheng2024gpt, mobile agents are embedded in smartphone applications to automate mobile tasks zhang2025appagent; wang2024mobile, and desktop or laptop agents interact with local operating systems and graphical user interfaces zhang2025ufo; zhang2025ufo2. Across these platforms, agents observe system state, reason about tasks, and execute actions, enabling applications such as automated customer support flows, desktop productivity macros, and repetitive workflow automation, essentially acting as intelligent digital assistants.
Despite their versatility and growing adoption, existing digital agents are fundamentally limited by their confinement to a single device or environment. In practice, many modern workflows are no longer confined to a single endpoint brudy2018investigating: users frequently interact with a combination of desktops, mobile devices, cloud services, and specialized hardware, all as part of a single task xu2021cross; chen2020multi. Examples include running data analysis on a GPU cluster, collecting results on a personal laptop, and generating visual summaries on a tablet, or coordinating cross-platform deployments that touch both local workstations and cloud infrastructure. This trend makes the ability to operate seamlessly across devices increasingly urgent and commonplace.
@@
  • Fragmented personalization and context. User-specific preferences, personalization, and contextual knowledge captured on one device are difficult to transfer or leverage on another cemri2025multi, preventing a coherent multi-device experience.
@@ 2.2 Cross-Device Agent: A New Paradigm
To overcome the limitations of single-device frameworks, we introduce the concept of cross-device intelligence, in which digital agents collaborate seamlessly across multiple heterogeneous endpoints brudy2019cross. In this metaphorical digital galaxy, each device, desktop, mobile, cloud service, or specialized hardware, acts as a star, and coordinated tasks form a constellation that collectively fulfills complex user requests. Unlike traditional agents confined to a single host, these cross-device agents can reason about device capabilities, orchestrate subtasks, and execute actions as part of a unified, intelligent ecosystem cheninternet
@@
This paradigm directly addresses the key limitations of single-device agents. It overcomes the constraints of individual devices by assigning tasks to the endpoints with the necessary capabilities and resources. It allows personalization, context, and user preferences captured on one device to propagate across others, enabling a coherent multi-device experience giusti2025federation. And it automates orchestration of complex workflows, removing reliance on brittle glue code or labor-intensive manual integration.
@@ 2.3 Design Challenges in Cross-Device Orchestration
However, synchronous control loops, cross-device agent orchestration introduces a fundamentally different class of systems challenges. In a distributed, heterogeneous environment, agents must collaborate across device, network, and platform boundaries, each with distinct runtime contexts and execution semantics tran2025multi. Building a reliable and efficient orchestration layer under such conditions requires addressing three core challenges.
@@
    Asynchronous Parallelism. In cross-device workflows, multiple agents may execute concurrently on different endpoints. Unlike linear single-agent plans, task execution must accommodate partial completions, delayed feedback, and dynamic dependency resolution. The orchestrator must therefore reason about concurrency, detect when subtasks can safely proceed in parallel, and continuously adapt its scheduling plan based on evolving runtime states and network latencies yu2025dyntaskmas. Failing to do so may lead to wasted computation or stalled progress.
@@
    Distributed Coordination. Agents in a constellation operate across diverse network and trust domains, implemented in different languages and deployed on various infrastructures. Achieving coherent coordination among them requires standardized mechanisms for registration, capability advertisement, task dispatch, and result collection. These operations must occur through persistent, low-latency channels that can tolerate temporary disconnections and guarantee consistent task states cheninternet. In practice, ad-hoc HTTP calls or ephemeral connections are insufficient; a structured communication substrate is essential to sustain long-lived agent interactions.
@@
    Heterogeneous Extensibility. A cross-device ecosystem must embrace diversity rather than constrain it. Device agents differ in operating systems, execution environments, and available toolchains. The architecture should therefore allow rapid development, deployment, and integration of new device agents while ensuring consistent semantics in task execution and error propagation wu2024autogen. This requires a flexible interface model and a unified protocol to abstract platform-specific complexity, enabling scalable and evolvable orchestration across an open agent federation yang2025survey.
@@ 4.4 Runtime Dynamism and Adaptivity
Unlike static DAG schedulers Islam2012Oozie; di2017nextflow; argo, UFO3 treats TaskConstellations as mutable objects. Tasks and dependency edges can be inserted, removed, or modified at runtime. This design enables UFO3 to dynamically react to evolving execution contexts, such as new user inputs, intermediate results, or device failures, without restarting the entire workflow. Such adaptivity is key to maintaining progress and efficiency in long-running, cross-device task orchestration.
@@ 5 ConstellationAgent: The Centralized Constellation Weaver
As illustrated in Figure 4, ConstellationAgent acts as both a planner and a replanner: it first synthesizes a TaskConstellation from user instructions, then incrementally refines this constellation as feedback arrives from distributed agents. Internally, ConstellationAgent is implemented as an LLM-driven ReAct agent yao2022react governed by a finite-state machine (FSM) schneider2005state. This FSM alternates between two complementary operating modes, namely creation and editing, forming a closed-loop control cycle that continuously updates the global execution graph in response to runtime conditions.
@@ 5.1 Responsibilities and I/O Interface
In the creation phase, ConstellationAgent receives three primary inputs: (i) a user-issued goal, expressed in natural or structured language; (ii) the AgentProfile registry, describing each available device agent’s capabilities, environment, and metadata; and (iii) demonstration examples to support in-context learning (ICL) dong2022survey; jiang2024xpert.
@@
The TaskConstellation is generated in a structured, machine-readable JSON format. Beyond the raw graph, ConstellationAgent produces a detailed reasoning trace: it first formulates an Observation of the input and AgentProfile, followed by a structured Thought capturing its analysis and decision-making rationale wei2022chain; ding2024everything. It then outputs the next State in the control loop and the overall Result for user-request or the generated TaskConstellation. This ensures that the TaskConstellation is derived through a transparent, deliberate reasoning process.
@@ 5.3 Constellation MCP Server: Structured Task Management
To operationalize dynamic graph construction, ConstellationAgent interacts with a lightweight Constellation MCP Server hou2025model, a modular component exposing a standardized set of task and dependency management primitives. This server serves as the structured manipulation layer that bridges LLM-level reasoning and concrete execution state. Each operation encapsulates a single, idempotent transformation on the TaskConstellation, ensuring reproducibility and easy rollback.
@@ 6 Asynchronous Dynamic Constellation Orchestrator
With the ConstellationAgent governs the reasoning and evolution of a TaskConstellation, the Constellation Orchestrator brings that plan to life by executing, monitoring, and adapting interdependent tasks across heterogeneous devices li2022dag, as shown in Figure 6. Conceptually, it transforms a static DAG into a living execution fabric, where tasks evolve concurrently, react to runtime signals, and adapt to new decisions generated by the reasoning agent.
@@ 7 Agent Interaction Protocol (AIP)
The orchestration model described in Section 6 requires a communication substrate that remains correct under continuous DAG evolution, dynamic agent participation, and fine-grained event propagation. Legacy HTTP-based coordination approaches (e.g., A2A duan2025agent, ACP ehtesham2025survey) assume short-lived, stateless interactions, incurring handshake overhead, stale capability views, and fragile recovery when partial failures occur mid-task. These assumptions make them unsuitable for the continuously evolving workflows and long-running reasoning loops characteristic of UFO3.
@@ 10 Experimental Evaluation
Following the system implementation, we evaluate the UFO3 framework to understand its effectiveness, robustness, and coordination efficiency in realistic, heterogeneous environments. Existing agent benchmarks predominantly focus on single-device or single-OS settings (e.g., text-based API workflows or GUI automation) mu2025gui, which fail to capture the cross-device, cross-platform orchestration that UFO3 is designed for. Moreover, to the best of our knowledge, there exists no prior agent system capable of orchestrating multi-device tasks that span both Linux and Windows environments with unified control and shared context. This makes direct comparison against existing systems infeasible and potentially misleading.
@@ 10.1 Experiment Setup
Agents communicate via a simulated local network with 1–10 ms latency and a wide-area link (50–100 ms latency) for the GPU node using the AIP. All components are implemented in Python 3.10 and powered by the same large language model (GPT-5-Chat-20251003) openai2025gpt5systemcard for both the ConstellationAgent and device agents. The orchestrator and controller run on a dedicated management node that coordinates agent discovery, task scheduling, and monitoring. This setup emulates a realistic hybrid enterprise environment that combines cloud servers, local desktops, and GPU compute nodes.

### v2

# arXiv:2511.11332v2[cs.DC] 01 Mar 2026

@@ 1 Introduction
The rise of intelligent agents Wang et al. (2024b) marks a new era of human–computer interaction, where large language models (LLMs) Naveed et al. (2025) are evolving from text-based reasoning engines Qu et al. (2025) into autonomous digital operators capable of perceiving, acting, and coordinating across tasks Zhang et al. (2024). Yet despite this progress, most agent frameworks remain confined within a single device or platform, be it a browser tab Ning et al. (2025); Zheng et al. (2024), a desktop environment Zhang et al. (2025b; a), or a mobile app Zhang et al. (2025d); Wang et al. (2024a). This confinement sharply limits their ability to harness the rich, distributed computational ecosystem that modern users inhabit.
When an agent is trapped within one operating system, it cannot access the complementary strengths of other devices, such as GPU clusters for computation, desktop applications for document editing, or mobile sensors for context capture. The result is a fragmented landscape of intelligent but siloed agents: each powerful in isolation yet collectively underutilized. This gap between reasoning ability and real-world actuation leaves vast potential untapped. To truly advance the next frontier of automation and reasoning, agents must operate beyond the boundaries of any single device or OS, forming a coherent digital collective where Windows laptops, Linux servers, mobile devices, and edge nodes collaborate seamlessly Houben et al. (2017); Brudy et al. (2019) to gather and act upon ubiquitous intelligence.
@@
  • Declarative decomposition into a dynamic DAG (TaskConstellation). Natural-language or programmatic requests are decomposed by the global ConstellationAgent into a structured DAG Bei et al. (2025) of TaskStars and TaskStarLines that encode workflow logic and dependencies. This declarative structure is amenable to automated scheduling, introspection, and rewriting throughout execution.
  • Continuous, result-driven graph evolution. The TaskConstellation is a living data structure. Intermediate outputs, transient failures, and new observations trigger controlled rewrites, adding diagnostic TaskStars, creating fallbacks, rewiring dependencies, or pruning completed nodes, so the system adapts dynamically instead of aborting on errors Wu et al. (2024b).
@@
  • Unified Agent Interaction Protocol (AIP). Built atop persistent WebSocket channels, we develop AIP, a protocol that provides a unified, secure, and fault-tolerant layer for agent registry, session management, task dispatch, and coordination. It ensures reliability under network fluctuations through automatic reconnection and retry, while exposing a lightweight, extensible interface that allows new agents to integrate seamlessly into the UFO3 ecosystem Yang et al. (2025a).
  • Template-driven framework for MCP-empowered device agents. To democratize agent creation, UFO3 provides a lightweight development template and toolkit for rapidly building new device agents. Developers can declare capabilities, bind to local environments, and extend them through one or more Model Context Protocol (MCP) servers Hou et al. (2025) for tool augmentation. This modular design accelerates integration while maintaining consistency across the constellation.
@@
In essence, UFO3 dissolves device boundaries and transforms the digital estate into a single, adaptive collaborator Brudy et al. (2018); Marks and White (2020); Zhang et al. (2018). It unifies distributed devices into a cohesive digital organism, one that executes user intents safely, asynchronously, and efficiently across heterogeneous environments. At its core, the Agent Interaction Protocol (AIP) serves as the connective tissue of this ecosystem, roviding a unified, fault-tolerant, and extensible communication substrate that allows new agents to join seamlessly and interoperate reliably. Over time, multiple constellations can interconnect through AIP, weaving together agents, devices, and capabilities into a self-organizing Digital Agent Galaxy. Through this design, UFO3 redefines cross-device automation, elevating it from a brittle engineering challenge to a unified orchestration paradigm, where multi-device workflows become naturally expressive and scalable across the landscape of ubiquitous computing.
@@ 2.1 Digital Agents
Leveraging the power of LLMs, modern digital agents have emerged as powerful interfaces bridging human intent and the complex software ecosystems people interact with daily Zhang et al. (2024). These agents can parse natural language requests, interpret screenshots and system state Zheng et al. (2025); Wu et al. (2025); Zhao et al. (2025), decompose complex goals into subtasks Huang et al. (2024), and generate scripts or commands to execute tasks using a variety of tools Wang et al. (2024b). Execution modalities include API calls, operating-system-level commands, GUI interactions via automation or accessibility interfaces, and code generation Zhang et al. ; Qiao et al. (2023); Wang et al. (2024c); Zhang et al. (2025e).
Digital agents operate across a wide spectrum of platforms. Web-based agents can navigate browsers and search the Internet Ning et al. (2025); Zheng et al. (2024), mobile agents are embedded in smartphone applications to automate mobile tasks Zhang et al. (2025d); Wang et al. (2024a), and desktop or laptop agents interact with local operating systems and graphical user interfaces Zhang et al. (2025b; a). Across these platforms, agents observe system state, reason about tasks, and execute actions, enabling applications such as automated customer support flows, desktop productivity macros, and repetitive workflow automation, essentially acting as intelligent digital assistants.
Despite their versatility and growing adoption, existing digital agents are fundamentally limited by their confinement to a single device or environment. In practice, many modern workflows are no longer confined to a single endpoint Brudy et al. (2018): users frequently interact with a combination of desktops, mobile devices, cloud services, and specialized hardware, all as part of a single task Xu et al. (2021); Chen and Koufaris (2020). Examples include running data analysis on a GPU cluster, collecting results on a personal laptop, and generating visual summaries on a tablet, or coordinating cross-platform deployments that touch both local workstations and cloud infrastructure. This trend makes the ability to operate seamlessly across devices increasingly urgent and commonplace.
@@
  • Fragmented personalization and context. User-specific preferences, personalization, and contextual knowledge captured on one device are difficult to transfer or leverage on another Cemri et al. (2025), preventing a coherent multi-device experience.
@@ 2.2 Cross-Device Agent: A New Paradigm
To overcome the limitations of single-device frameworks, we introduce the concept of cross-device intelligence, in which digital agents collaborate seamlessly across multiple heterogeneous endpoints Brudy et al. (2019). In this metaphorical digital galaxy, each device, desktop, mobile, cloud service, or specialized hardware, acts as a star, and coordinated tasks form a constellation that collectively fulfills complex user requests. Unlike traditional agents confined to a single host, these cross-device agents can reason about device capabilities, orchestrate subtasks, and execute actions as part of a unified, intelligent ecosystem Chen et al.
@@
This paradigm directly addresses the key limitations of single-device agents. It overcomes the constraints of individual devices by assigning tasks to the endpoints with the necessary capabilities and resources. It allows personalization, context, and user preferences captured on one device to propagate across others, enabling a coherent multi-device experience Giusti et al. (2025). And it automates orchestration of complex workflows, removing reliance on brittle glue code or labor-intensive manual integration.
@@ 2.3 Design Challenges in Cross-Device Orchestration
However, synchronous control loops, cross-device agent orchestration introduces a fundamentally different class of systems challenges. In a distributed, heterogeneous environment, agents must collaborate across device, network, and platform boundaries, each with distinct runtime contexts and execution semantics Tran et al. (2025). Building a reliable and efficient orchestration layer under such conditions requires addressing three core challenges.
@@
    Asynchronous Parallelism. In cross-device workflows, multiple agents may execute concurrently on different endpoints. Unlike linear single-agent plans, task execution must accommodate partial completions, delayed feedback, and dynamic dependency resolution. The orchestrator must therefore reason about concurrency, detect when subtasks can safely proceed in parallel, and continuously adapt its scheduling plan based on evolving runtime states and network latencies Yu et al. (2025). Failing to do so may lead to wasted computation or stalled progress.
@@
    Distributed Coordination. Agents in a constellation operate across diverse network and trust domains, implemented in different languages and deployed on various infrastructures. Achieving coherent coordination among them requires standardized mechanisms for registration, capability advertisement, task dispatch, and result collection. These operations must occur through persistent, low-latency channels that can tolerate temporary disconnections and guarantee consistent task states Chen et al. . In practice, ad-hoc HTTP calls or ephemeral connections are insufficient; a structured communication substrate is essential to sustain long-lived agent interactions.
@@
    Heterogeneous Extensibility. A cross-device ecosystem must embrace diversity rather than constrain it. Device agents differ in operating systems, execution environments, and available toolchains. The architecture should therefore allow rapid development, deployment, and integration of new device agents while ensuring consistent semantics in task execution and error propagation Wu et al. (2024a). This requires a flexible interface model and a unified protocol to abstract platform-specific complexity, enabling scalable and evolvable orchestration across an open agent federation Yang et al. (2025a).
@@ 4.4 Runtime Dynamism and Adaptivity
Unlike static DAG schedulers M. Islam, A. K. Huang, M. Battisha, M. Chiang, S. Srinivasan, C. Peters, A. Neumann, and A. Abdelnur (2012); P. Di Tommaso, M. Chatzou, E. W. Floden, P. P. Barja, E. Palumbo, and C. Notredame (2017); 2, UFO3 treats TaskConstellations as mutable objects. Tasks and dependency edges can be inserted, removed, or modified at runtime. This design enables UFO3 to dynamically react to evolving execution contexts, such as new user inputs, intermediate results, or device failures, without restarting the entire workflow. Such adaptivity is key to maintaining progress and efficiency in long-running, cross-device task orchestration.
@@ 5 ConstellationAgent: The Centralized Constellation Weaver
As illustrated in Figure 4, ConstellationAgent acts as both a planner and a replanner: it first synthesizes a TaskConstellation from user instructions, then incrementally refines this constellation as feedback arrives from distributed agents. Internally, ConstellationAgent is implemented as an LLM-driven ReAct agent Yao et al. (2022) governed by a finite-state machine (FSM) Schneider (2005). This FSM alternates between two complementary operating modes, namely creation and editing, forming a closed-loop control cycle that continuously updates the global execution graph in response to runtime conditions.
@@ 5.1 Responsibilities and I/O Interface
In the creation phase, ConstellationAgent receives three primary inputs: (i) a user-issued goal, expressed in natural or structured language; (ii) the AgentProfile registry, describing each available device agent’s capabilities, environment, and metadata; and (iii) demonstration examples to support in-context learning (ICL) Dong et al. (2022); Jiang et al. (2024).
@@
The TaskConstellation is generated in a structured, machine-readable JSON format. Beyond the raw graph, ConstellationAgent produces a detailed reasoning trace: it first formulates an Observation of the input and AgentProfile, followed by a structured Thought capturing its analysis and decision-making rationale Wei et al. (2022); Ding et al. (2024). It then outputs the next State in the control loop and the overall Result for user-request or the generated TaskConstellation. This ensures that the TaskConstellation is derived through a transparent, deliberate reasoning process.
@@ 5.3 Constellation MCP Server: Structured Task Management
To operationalize dynamic graph construction, ConstellationAgent interacts with a lightweight Constellation MCP Server Hou et al. (2025), a modular component exposing a standardized set of task and dependency management primitives. This server serves as the structured manipulation layer that bridges LLM-level reasoning and concrete execution state. Each operation encapsulates a single, idempotent transformation on the TaskConstellation, ensuring reproducibility and easy rollback.
@@ 6 Asynchronous Dynamic Constellation Orchestrator
With the ConstellationAgent governs the reasoning and evolution of a TaskConstellation, the Constellation Orchestrator brings that plan to life by executing, monitoring, and adapting interdependent tasks across heterogeneous devices Li et al. (2022), as shown in Figure 6. Conceptually, it transforms a static DAG into a living execution fabric, where tasks evolve concurrently, react to runtime signals, and adapt to new decisions generated by the reasoning agent.
@@ 7 Agent Interaction Protocol (AIP)
The orchestration model described in Section 6 requires a communication substrate that remains correct under continuous DAG evolution, dynamic agent participation, and fine-grained event propagation. Legacy HTTP-based coordination approaches (e.g., A2A Duan and Lu (2025), ACP Ehtesham et al. (2025)) assume short-lived, stateless interactions, incurring handshake overhead, stale capability views, and fragile recovery when partial failures occur mid-task. These assumptions make them unsuitable for the continuously evolving workflows and long-running reasoning loops characteristic of UFO3.
@@ 10 Experimental Evaluation
Following the system implementation, we evaluate the UFO3 framework to understand its effectiveness, robustness, and coordination efficiency in realistic, heterogeneous environments. Existing agent benchmarks predominantly focus on single-device or single-OS settings (e.g., text-based API workflows or GUI automation) Mu et al. (2025), which fail to capture the cross-device, cross-platform orchestration that UFO3 is designed for. Moreover, to the best of our knowledge, there exists no prior agent system capable of orchestrating multi-device tasks that span both Linux and Windows environments with unified control and shared context. This makes direct comparison against existing systems infeasible and potentially misleading.
@@ 10.1 Experiment Setup
Agents communicate via a simulated local network with 1–10 ms latency and a wide-area link (50–100 ms latency) for the GPU node using the AIP. All components are implemented in Python 3.10 and powered by the same large language model (GPT-5-Chat-20251003) OpenAI (2025b) for both the ConstellationAgent and device agents. The orchestrator and controller run on a dedicated management node that coordinates agent discovery, task scheduling, and monitoring. This setup emulates a realistic hybrid enterprise environment that combines cloud servers, local desktops, and GPU compute nodes.